# Dynamic Risk-Aware MPPI for Mobile Robots in Crowds via Efficient Monte Carlo Approximations

### v1

@@ Abstract
Deploying mobile robots safely among humans requires the motion planner to account for the uncertainty in the other agents’ predicted trajectories. This remains challenging in traditional approaches, especially with arbitrarily shaped predictions and real-time constraints. To address these challenges, we propose a Dynamic Risk-Aware Model Predictive Path Integral control (DRA-MPPI), a motion planner that incorporates uncertain future motions modelled with potentially non-Gaussian stochastic predictions. By leveraging MPPI’s gradient-free nature, we propose a method that efficiently approximates the joint Collision Probability (CP) among multiple dynamic obstacles for several hundred sampled trajectories in real-time via a Monte Carlo (MC) approach. This enables the rejection of samples exceeding a predefined CP threshold or the integration of CP as a weighted objective within the navigation cost function. Consequently, DRA-MPPI mitigates the freezing robot problem while enhancing safety. Real-world and simulated experiments with multiple dynamic obstacles demonstrate DRA-MPPI’s superior performance compared to state-of-the-art approaches, including Scenario-based Model Predictive Control (S-MPC), Frenét planner, and vanilla MPPI.
@@ I Introduction
Mobile robots have the potential to enhance various aspects of daily life, from optimizing logistics in warehouses to enabling safer and more efficient transportation through autonomous vehicles [1, 2, 3]. However, many existing robotic systems operate in structured environments with minimal human interaction, such as designated lanes or controlled spaces, to simplify navigation. For robots to be successfully integrated into real-world settings like urban areas, they must be capable of safely and efficiently manoeuvring through human-populated spaces. Achieving this requires an ability to interpret and anticipate human movement—a task complicated by the inherent unpredictability of human behaviour. Prediction models, such as [4], provide probabilistic distributions over potential human trajectories. To ensure safe and efficient navigation, these probabilistic predictions must be incorporated into the motion planning process [5, 6, 7].
@@ III-A Dynamic Obstacle Model
The motion planner relies on a predictive model that probabilistically forecasts the behaviour of surrounding agents. In this setting, we model the positions of $N_{o}$ dynamic obstacles as random variables. Specifically, the uncertain position of obstacle $o$ at time step $t$ as $\bm{\delta}^{o}_{t}=[x_{t}^{o},y_{t}^{o}]\in\Delta$, where $\Delta$ represents the probability space capturing the uncertainty in the perception of these obstacles. This uncertainty is formally characterized by a probability measure $\mathbb{P}$. To capture the inherent multi-modality in obstacle behaviour, we model the uncertainty in their motion using a Mixture-of-Gaussians (MoG) distribution. This approach provides a flexible and expressive representation of uncertainty by combining multiple continuous probability distributions. Specifically, the probability density function of an obstacle’s position at time step $t$ is given by:
@@ III-B Probabilistic Collision Avoidance
$\mathbb{P}\left[||\bm{q}_{t}-\bm{\delta}_{t}^{o}||_{2}>r,\forall o\right]\geq 1-\sigma,\forall t,$ (4d)
@@ III-C Model Predictive Path Integral Control (MPPI)
Model Predictive Path Integral control (MPPI) [21] is a sampling-based optimization method for solving stochastic optimal control problems. We consider a discrete-time dynamical system governed by the state-transition equation:
@@ III-D Risk Assessment
[…]
@@ IV-A Monte Carlo Approximation
$\hat{\mathcal{P}}(x_{k,t},y_{k,t})=\frac{\pi r^{2}}{N_{\mathcal{R}^{\circ}_{k,t}}}\sum_{j=1}^{N_{mc}}\mathcal{P}^{\text{joint}}(x_{j},y_{j})\cdot\mathbbm{1}_{\mathcal{R}^{\circ}_{k,t}}(x_{j},y_{j}),$ (12)
where $\pi r^{2}$ is the area of the collision region and $N_{\mathcal{R}^{\circ}_{k,t}}$ is the number of Monte Carlo samples in the circular region $\mathcal{R}^{\circ}_{k,t}$ acting as a normalization factor. $\mathbbm{1}_{\mathcal{R}^{\circ}_{k,t}}$ is an indicator function to consider only MC samples in the circular collision region $\mathcal{R}^{\circ}_{k,t}$. Given that all the coordinates $x_{k,t}$ and $y_{k,t}$ will be fairly close $\forall k$, there will also be significant overlap in the integration regions $\mathcal{R}^{\circ}_{k,t}$. Thus, the computation effort is massively reduced by using our approach of drawing and evaluating the MC samples only once $\forall k$. The Standard Error (SE) of $\hat{\mathcal{P}}$, if $\mathcal{P}^{\text{joint}}$ were Gaussian distributed, would scale with $\sqrt{\Sigma_{\mathcal{P}^{\text{joint}}}}/N_{\mathcal{R}^{\circ}_{k,t}}$. In practice, we do not know a priori the number of samples in the collision region, and $\mathcal{P}^{\text{joint}}$ is non-gaussian but, compared to an analytical solution, we have empirically observed that we misclassify the CP to be below the threshold less than 2% of the times.
@@ IV-B Cost Formulation
$\mathcal{C}_{\text{risk}}=\omega_{\text{soft}}\hat{\mathcal{P}}(\bm{q}_{k,t})+\omega_{\text{hard}}\mathbbm{1}_{\sigma}(\hat{\mathcal{P}}(\bm{q}_{k,t})),$ (14)
with $\omega_{\text{soft}}$ and $\omega_{\text{hard}}$ being tunable weights. The first part, $\omega_{\text{soft}}\hat{\mathcal{P}}(\bm{q}_{k,t})$, is a linear penalty on any collision probabilities, incentivising the planner to always choose lower risk. The second part, $\omega_{\text{hard}}\mathbbm{1}_{\sigma}(\hat{\mathcal{P}}(\bm{q}_{k,t}))$ is an indicator function returning $1$ when the CP is above a chosen risk threshold $\sigma$. Every sampled trajectory with a CP above the threshold, assuming that $\omega_{\text{hard}}$ is high enough and that at least a few low-risk trajectories are sampled, will receive close to zero weight in the average in 7, essentially being rejected. One advantage is that if, for example, the robot starts planning from a state with CP above the threshold, MPPI still returns a solution in which the CP and the number of time steps with CP above the threshold will be minimised.
@@ IV-C Algorithm
2:      $U_{init}=[0,\dots,0]$triangleright $U_{init}\in\mathbb{R}^{T}$
[…]
@@
9:              $\mathcal{R}^{{\scriptscriptstyle\square}}_{t}\leftarrow\textit{getArea}(\left[\bm{x}_{0,t},...,\bm{x}_{K,t}\right])$triangleright 10
@@
12:                  $\mathcal{P}^{\text{joint}}_{j}\leftarrow\textit{evaluateCP}(x_{j},y_{j})$triangleright 11
@@
14:              $\hat{\mathcal{P}}_{k,t+1}\leftarrow&\textit{estimateJointCP}(\bm{x}_{k,t+1},\\ &\left[\mathcal{P}^{\text{joint}}_{1},...,\mathcal{P}^{\text{joint}}_{N_{mc}}\right])\end{aligned}$triangleright 12
@@
16:         $S_{k}\leftarrow\textit{getCost}({X}_{k},\left[\hat{\mathcal{P}}_{k,1},...,\hat{\mathcal{P}}_{k,T}\right])$triangleright 13
[…]
@@ V-A Simulation Setup
The experimental evaluation is conducted in a simulated environment featuring a Clearpath Jackal robot navigating a 6-meter-wide corridor alongside multiple pedestrians. The robot has to follow the corridor’s centerline while maintaining a reference velocity of $2$m/s. A visualization with Gaussian predictions is given in Fig. 1. The simulator runs at $20$Hz, while the controllers run at $5$Hz. Pedestrian dynamics are modelled using the Social Forces [28], implemented via open-source software [29]. In this model, pedestrians interact with one another and the robot. Future pedestrian positions are predicted by the robot’s motion planner using a constant velocity assumption. Pedestrians are modelled with a radius of $0.3$m and spawn on both sides of the corridor, and their goal is to reach the opposing side. The robot dynamics are described by a second-order unicycle model [30]. Simulations are conducted for three scenarios featuring $4$, $8$, and $12$ pedestrians to evaluate the planner’s performance under varying crowd density levels.
@@ V-A1 Pedestrians with Gaussian noise
In this setup, we model the uncertainty in pedestrian motion using an unimodal Gaussian distribution with covariance matrix $\bm{\Sigma}_{w}=0.3^{2}\bm{I}$. We define the pedestrians’ dynamics as:
@@ V-A2 Pedestrians with Mixture of Gaussians
# Pedestrians Method Task Duration [s] Velocity [m/s] CP Safe (%) Runtime [ms]
@@
# Pedestrians Method Task Duration [s] Velocity [m/s] CP Safe (%) Runtime [ms]
@@ V-B Comparison to Baselines
    Safety: Evaluates the proportion of experiments in which the robot successfully avoids collisions with pedestrians.
@@
Table I and Fig. 2 and 3 present a quantitative comparison of different planning methods across varying pedestrian densities under uni-modal pedestrian predictions. The results, averaged over $100$ experiments, highlight the trade-offs between task efficiency and safety among the evaluated approaches. Our proposed approach (DRA-MPPI) consistently demonstrates a favourable balance between task duration and safety. While Vanilla MPPI achieves the shortest task duration across all scenarios, it does so at the expense of significantly lower safety due to its lack of risk awareness, particularly as pedestrian density increases. For instance, in the presence of $12$ pedestrians, Vanilla MPPI completes the task in $18.98$s but exhibits the lowest safety rate ($43$%), indicating a high collision frequency. In contrast, DRA-MPPI maintains a competitive task duration ($19.86$s) while achieving a significantly higher safety rate of $98$%, demonstrating its robustness in crowded environments. The Frenét Planner exhibits longer task duration and suffers a notable drop in safety, particularly at higher pedestrian densities ($74\%$ safety at 12 pedestrians). This deterioration arises from the planner’s design, which does not explicitly optimize for risk minimization. Instead, it evaluates a set of sampled trajectories based on assigned costs and selects the one with the lowest one. However, in dense environments, the trajectory with the minimum cost may still entail substantial risk, as observed in the experimental results.
@@
SH-MPC strives to track the reference path very closely, often finding itself in riskier positions. SH-MPC relaxes the constraints with slack variables when no feasible solution is found in these riskier situations. When that does not work, SH-MPC may find no solution and return zero input (braking). This leads to SH-MPC taking longer to complete the task and often exhibiting collision probabilities higher than the threshold at least a few times per experiment, severely impacting the metrics.
@@ V-C Computation Time
In Tables I and II, we also report the computation time (runtime) for SH-MPC and our approach. DRA-MPPI takes just below $100$ms to compute the control action in the uni-modal four-agents experiment, while SH-MPC is almost twice as fast. Both methods scale quite gracefully over the number of agents with multi-modal predictions. However, it is important to note that SH-MPC is written in C++, while DRA-MPPI is implemented in PyTorch. Recent work shows a reduction of compute time of up to two orders of magnitude when running MPPI in CUDA vs Pytorch [33], highlighting the potential for these massively parallelizable approaches.
@@ VII Conclusions
Real-robot experiments showcased DRA-MPPI’s ease of transfer to the real world and real-time performance.

### v2

@@ Abstract
Deploying mobile robots safely among humans requires the motion planner to account for the uncertainty in the other agents’ predicted trajectories. This remains challenging in traditional approaches, especially with arbitrarily shaped predictions and real-time constraints. To address these challenges, we propose a Dynamic Risk-Aware Model Predictive Path Integral control (DRA-MPPI), a motion planner that incorporates uncertain future motions modelled with potentially non-Gaussian stochastic predictions. By leveraging MPPI’s gradient-free nature, we propose a method that efficiently approximates the joint Collision Probability (CP) among multiple dynamic obstacles for several hundred sampled trajectories in real-time via a Monte Carlo (MC) approach. This enables the rejection of samples exceeding a predefined CP threshold or the integration of CP as a weighted objective within the navigation cost function. Consequently, DRA-MPPI mitigates the freezing robot problem while enhancing safety. Real-world and simulated experiments with multiple dynamic obstacles demonstrate DRA-MPPI’s superior performance compared to state-of-the-art approaches, including Scenario-based Model Predictive Control (S-MPC), Frenét planner, and vanilla MPPI. Videos of the experiments can be found at https://autonomousrobots.nl/paper_websites/dra-mppi.
@@ I Introduction
Mobile robots have the potential to enhance various aspects of daily life, from optimizing logistics in warehouses to enabling safer and more efficient transportation through autonomous vehicles [1, 2, 3]. However, for robots to be successfully integrated into real-world settings like urban areas, they must be capable of safely and efficiently manoeuvring through human-populated spaces. Achieving this requires an ability to interpret and anticipate human movement—a task complicated by the inherent unpredictability of human behaviour. Prediction models, such as [4], provide probabilistic distributions over potential human trajectories. To ensure safe and efficient navigation, these probabilistic predictions must be incorporated into the motion planning process [5, 6, 7].
@@ III-A Dynamic Obstacle Model
The motion planner uses a model that probabilistically forecasts the behaviour of surrounding agents. In this setting, we model the positions of $N_{o}$ dynamic obstacles as random variables. Specifically, the uncertain position of obstacle $o$ at time step $t$ as $\bm{\delta}^{o}_{t}=[x_{t}^{o},y_{t}^{o}]\in\Delta$, where $\Delta$ is the probability space capturing the uncertainty in the perception of these obstacles. This uncertainty is characterized by a probability measure $\mathbb{P}$. To capture the inherent multi-modality in obstacle behaviour, we model the uncertainty in their motion using a Mixture-of-Gaussians (MoG) distribution. This approach provides a flexible and expressive representation of uncertainty by combining multiple continuous probability distributions. Specifically, the probability density function of an obstacle’s position at time step $t$ is given by:
@@ III-B Probabilistic Collision Avoidance
[…]
@@ III-C Model Predictive Path Integral Control (MPPI)
Model Predictive Path Integral control (MPPI) [21, 24] is a sampling-based optimization method for solving stochastic optimal control problems. We consider a discrete-time dynamical system governed by the state-transition equation:
@@ III-D Risk Assessment
$\mathcal{P}_{t}^{o}(\bm{q}_{t})=\iint_{x_{t},y_{t}\in\mathcal{R}^{\circ}_{t}}p_{t}^{o}(x_{t},y_{t})dxdy,\forall{t},o,$ (8)
@@ IV-A Monte Carlo Approximation
$\hat{\mathcal{P}}(x_{k,t},y_{k,t})=\frac{\pi r^{2}}{N_{\mathcal{R}^{\circ}_{k,t}}}\sum_{j=1}^{N_{mc}}\mathcal{P}^{\text{joint}}(x_{j},y_{j})\cdot\mathds{1}_{\mathcal{R}^{\circ}_{k,t}}(x_{j},y_{j}),$ (12)
where $\pi r^{2}$ is the area of the collision region and $N_{\mathcal{R}^{\circ}_{k,t}}$ is the number of Monte Carlo samples in the circular region $\mathcal{R}^{\circ}_{k,t}$ acting as a normalization factor. $\mathds{1}_{\mathcal{R}^{\circ}_{k,t}}$ is an indicator function to consider only MC samples in the circular collision region $\mathcal{R}^{\circ}_{k,t}$. Given that all the coordinates $x_{k,t}$ and $y_{k,t}$ will be fairly close $\forall k$, there will also be significant overlap in the integration regions $\mathcal{R}^{\circ}_{k,t}$. Thus, the computation effort is massively reduced by using our approach of drawing and evaluating the MC samples only once $\forall k$. The Standard Error (SE) of $\hat{\mathcal{P}}$, if $\mathcal{P}^{\text{joint}}$ were Gaussian distributed, would scale with $\sqrt{\Sigma_{\mathcal{P}^{\text{joint}}}}/N_{\mathcal{R}^{\circ}_{k,t}}$. In practice, we do not know a priori the number of samples in the collision region, and $\mathcal{P}^{\text{joint}}$ is non-gaussian but, compared to an analytical solution, we have empirically observed that we misclassify the CP to be below the threshold less than 2% of the times.
@@ IV-B Cost Formulation
$\mathcal{C}_{\text{risk}}=\omega_{\text{soft}}\hat{\mathcal{P}}(\bm{q}_{k,t})+\omega_{\text{hard}}\mathds{1}_{\sigma}(\hat{\mathcal{P}}(\bm{q}_{k,t})),$ (14)
with $\omega_{\text{soft}}$ and $\omega_{\text{hard}}$ being tunable weights. The first part, $\omega_{\text{soft}}\hat{\mathcal{P}}(\bm{q}_{k,t})$, is a linear penalty on any collision probabilities, incentivising the planner to always choose lower risk. The second part, $\omega_{\text{hard}}\mathds{1}_{\sigma}(\hat{\mathcal{P}}(\bm{q}_{k,t}))$ is an indicator function returning $1$ when the CP is above a chosen risk threshold $\sigma$. Every sampled trajectory with a CP above the threshold, assuming that $\omega_{\text{hard}}$ is high enough and that at least a few low-risk trajectories are sampled, will receive close to zero weight in the average in 7, essentially being rejected. One advantage is that if, for example, the robot starts planning from a state with CP above the threshold, MPPI still returns a solution in which the CP and the number of time steps with CP above the threshold will be minimised.
@@ IV-C Algorithm
2:   $U_{init}=[0,\dots,0]$ $\triangleright$ $U_{init}\in\mathbb{R}^{T}$
$\mathcal{E}_{k}\leftarrow\textit{sampleHaltonSplines}()$ $\triangleright$ k=1Kk=1...K
@@
9:     $\mathcal{R}^{{\scriptscriptstyle\square}}_{t}\leftarrow\textit{getArea}(\left[\bm{x}_{0,t},...,\bm{x}_{K,t}\right])$ $\triangleright$ 10
@@
12:      $\mathcal{P}^{\text{joint}}_{j}\leftarrow\textit{evaluateCP}(x_{j},y_{j})$ $\triangleright$ 11
@@
14:     $\begin{aligned} \hat{\mathcal{P}}_{k,t+1}\leftarrow&\textit{estimateJointCP}(\bm{x}_{k,t+1},\\ &\left[\mathcal{P}^{\text{joint}}_{1},...,\mathcal{P}^{\text{joint}}_{N_{mc}}\right])\end{aligned}$ $\triangleright$ 12
@@
16:   $S_{k}\leftarrow\textit{getCost}({X}_{k},\left[\hat{\mathcal{P}}_{k,1},...,\hat{\mathcal{P}}_{k,T}\right])$ $\triangleright$ 13
17:   $w_{k}\leftarrow\textit{importanceSampling}$ $\triangleright$ 6
@@ V-A Simulation Setup
The experimental evaluation is conducted in a simulated environment featuring a Clearpath Jackal robot navigating a 6-meter-wide corridor alongside multiple pedestrians. The robot has to follow the corridor’s centerline while maintaining a reference velocity of 22m/s. A visualization with Gaussian predictions is given in Fig. 1. The simulator runs at 2020Hz, while the controllers run at 55Hz. Pedestrian dynamics are modelled using the Social Forces [29], implemented via open-source software [30]. In this model, pedestrians avoid one another and the robot. Future pedestrian positions are predicted by the robot’s motion planner using a constant velocity assumption. The mismatch between the model used for simulation and prediction is to emulate a real-world scenario where the predictions do not always match the ground-truth trajectory. Pedestrians are modelled with a radius of 0.30.3m and spawn on both sides of the corridor, and their goal is to reach the opposing side. The robot dynamics are described by a second-order unicycle model [31]. Simulations are conducted for three scenarios featuring $4$, 88, and $12$ pedestrians to evaluate the planner’s performance under varying crowd density levels.
@@ V-A1 Pedestrians with Gaussian noise
In this setup, we model the uncertainty in pedestrian motion using a unimodal Gaussian distribution with covariance matrix $\bm{\Sigma}_{w}=0.3^{2}\bm{I}$. By ‘unimodal’, it is meant that we predict a single trajectory for each pedestrian. We define the pedestrians’ dynamics as:
@@ V-A2 Pedestrians with Mixture of Gaussians
# Pedestrians Method Task Duration [s] Velocity [m/s] CP SR %) Runtime [ms]
@@
# Pedestrians Method Task Duration [s] Velocity [m/s] CP SR (%) Runtime [ms]
@@ V-B Comparison to Baselines
    Success Rate (SR): Evaluates the proportion of experiments in which the robot successfully avoids collisions with pedestrians and reaches its goal.
@@
Table I and Fig. 2 and 3 present a quantitative comparison of different planning methods across varying pedestrian densities under uni-modal pedestrian predictions. The results, averaged over $100$ experiments, highlight the trade-offs between task efficiency and safety among the evaluated approaches. Our proposed approach (DRA-MPPI) consistently demonstrates a favourable balance between task duration and safety. While Vanilla MPPI achieves the shortest task duration across all scenarios, it does so at the expense of significantly lower safety due to its lack of risk awareness, particularly as pedestrian density increases. For instance, in the presence of $12$ pedestrians, Vanilla MPPI completes the task in $18.98$s but exhibits the lowest safety rate (4343%), indicating a high collision frequency. In contrast, DRA-MPPI maintains a competitive task duration ($19.86$s) while achieving a significantly higher safety rate of 9898%, demonstrating its robustness in crowded environments. The Frenét Planner exhibits longer task duration and suffers a notable drop in safety, particularly at higher pedestrian densities ($74\%$ safety at 12 pedestrians). This deterioration arises from the planner’s design, which does not explicitly optimize for risk minimization. Instead, it evaluates a set of sampled trajectories based on assigned costs and selects the one with the lowest one. However, in dense environments, the trajectory with the minimum cost may still entail substantial risk, as observed in the experimental results.
@@
SH-MPC strives to track the reference path very closely, often finding itself in riskier positions. SH-MPC relaxes the constraints with slack variables when no feasible solution is found in these riskier situations. When infeasible, SH-MPC switches to a fallback policy (braking). This leads to SH-MPC taking longer to complete the task and often exhibiting collision probabilities higher than the threshold at least a few times per experiment, severely impacting the metrics.
@@ V-C Computation Time
In Tables I and II, we also report the computation time (runtime) for SH-MPC and our approach. DRA-MPPI takes just below 100100ms to compute the control action in the uni-modal four-agents experiment, while SH-MPC is almost twice as fast. Both methods scale quite gracefully over the number of agents with multi-modal predictions. However, it is important to note that SH-MPC is written in C++, while DRA-MPPI is implemented in PyTorch. Recent work shows a reduction of compute time of up to two orders of magnitude when running MPPI in CUDA vs Pytorch [34], highlighting the potential for these massively parallelizable approaches.
@@ VII Conclusions
Real-robot experiments showcased DRA-MPPI’s ease of transfer to the real world and real-time performance. Although we relied on a motion capture system, the pedestrian’s state and the map can be retrieved from onboard sensors [35].